OpenStudy (anonymous):

The number of farms in a certain area can be represented by the equation, y = 85,000(0.95)x, where x represents the number of years after 1990. What was the number of farms in 1990 (x = 0)? A. 95,000 farms B. 85,000 farms C. 80,750 farms D. 100,000 farms

OpenStudy (mathmale):

This problem might be clearer for you, as well as for others, if you'd please use the Equation Editor to write y = 85,000(0.95)x properly. (0.95)x normally means "0.95 times x," which is not correct here.\[y = 85,000(0.95)x \rightarrow y=85000(0.95) ^{x}\] is much clearer and less likely to be misinterpreted. Now, please substitute 0 for x in the above formula. What is the value of\[(0.95)^0?\]

OpenStudy (mathmale):

Continue on to find the value of y when x=0. Good luck!

OpenStudy (anonymous):

c ?

OpenStudy (mathmale):

T: Please show your work. I respond to explanations of your reasoning, or to images of your written work, but won't be telling you "right" or "wrong."

OpenStudy (anonymous):

i multiplied 85,000(0.95) and got 80,750